# Causal Explanations for Image Classifiers

### v2

@@ 5.2 Iterative refinement of responsibility
OUTPUT:   a responsibility map $\mathds{P}_{i}\longrightarrow\mathds{Q}$
@@
[…]
@@
One partition is rarely sufficient for a high-quality explanation, therefore we compute Algorithm 1 many times and compose the results, as shown in Algorithm 2. This calculates the responsibility for each superpixel (Line 1). If the termination condition is met (Lines 2–3), the responsibility map $Q$ is updated accordingly. Otherwise, for each superpixel in $\mathds{P}_{i}$, we refine it and call the algorithm recursively. We use $\cup$ to include these newly computed values in the returned map. The algorithm terminates when: 1) the superpixels in $\mathds{P}_{i}$ are sufficiently refined (containing only very few pixels), or 2) when all superpixels in $\mathds{P}_{i}$ have the same responsibility (this condition is for efficiency).
5.3 Explanation extraction
So far, we assume one particular partition $\mathds{P}_{i}$, which Algorithm 2 recursively refines and calculates the corresponding responsibilities of superpixels in each step by calling Algorithm 1. We note that the choice of the initial partition over the image can affect the values calculated by Algorithm 2, as this partition determines the set of possible mutants in Algorithm 1. We ameliorate the influence of the choice of any particular partition by iterating the algorithm over a set of partitions. Twenty iterations of the algorithm will therefore yield 202020 starting partitions, chosen at random. rex allows the user to choose from a number of different types of random to build these partitions, with uniform being the default.
@@ 5.3 Explanation extraction
[…]
@@
[…]
@@
The remainder of Algorithm 3 follows the method for explaining the result of an image classifier by (?). That is, we construct a subset of pixels $\mathcal{E}$ to explain $\mathcal{N}$’s output on this particular input $x$ greedily. We add pixels to $\mathcal{E}$ as long as $\mathcal{N}$’s output on $\mathcal{E}$ does not match $\mathcal{N}(x)$. This process terminates when $\mathcal{N}$’s output is the same as on the whole image $x$. The set $\mathcal{E}$ is returned as an explanation. We note that this extraction procedure is not normally followed by other familiar XAI tools, as their several different definitions of explanation do not extend to finding minimal sets of pixels which induce the desired classification.
While we approximate the computation of an explanation in order to ensure efficiency of our approach, the algorithm is built on solid theoretical foundations, which distinguishes it from other random or heuristic-based approaches. In practice, while our algorithm uses an iterative average of a greedy approximation, it yields highly accurate results (Section 6). Furthermore, our approach is simple and general, and uses the neural network as a blackbox.
5.4 Illustrative example
@@
Figure 3: The rex algorithm in action: rex creates an initial random partition of an image into 444 sections (①). All combinations of these sections are queried by the model, with further random refinement applied to those sections or combinations of sections that meet the requirements (②). Some combinations, highlighted in green, are classified as bus, others, in red, are not. This results, after several iterations, in a detailed responsibility map (③), from which a minimal passing explanations can be extracted (④).
To illustrate how rex works, consider Figure 3, which is classified as 'bus' by ResNet50, even though there is an occlusion in the middle. Initially, rex picks up an arbitrary partition of the image with four superpixels. This results in 151515 combinations of masking superpixels, though in practice, we only need to examine 141414 combinations as the $15^{th}$ is the entire image, which we have already tested. We use Algorithm 1 to calculate the responsibility of each superpixel. Those superpixels or combinations of superpixels which contribute towards the correct classification of further refined. This iterative refinement is a recursive application of the same initial random partitioning into four superpixels, as in Algorithm 2.
The responsibility mapss shown in Figure 4 demonstrate the importance of Algorithm 3. The initial partitioning of the image can greatly affect the quality of the responsibility map. Figure 4(a) shows the responsibility map after one iteration of the algorithm. While it still indicates well the areas of interest in the image it is also quite rectilinear. Figure 4(b) contains a lot more information about the distribution of responsibility over the pixels in the image. We now have two distinct peaks of responsibility on either side of the central occlusion (a person). Further iterations cause further refinement. By iteration 100 (Figure 4(e)), it is clear that the dominant explanation is around the rear tyre of the bus. The other explanation however, while depressed in relation to the tyre, has not vanished and can still be found through searching the responsibility map (?). Indeed, a disjoint explanation exists for this image, as given in Figure 5(a).
@@
Figure 4: Improvement of rex’s pixel ranking as the number of iterations $N$ increases (Algorithm 3)
@@
Figure 5: Two different explanations extracted from the 303030 iteration ranking in Figure 4Figure 5(a) contrasts interestingly with Figure 5(b): it seems that the front tire by itself is not sufficient (a very small rectangular section of the rear tire is included in the explanation), but the rear tire is.
5.5 Termination and complexity
It is clear that the algorithm terminates as soon as the parts can no longer be divided into superpixels. Moreover, the number of calls it performs to the model is linear in the size of the image, as proved in the following lemma.
@@ 5.5 Termination and complexity
The computation of responsibilities of superpixels in one partition is $O(2^{s})$, as the algorithm examines the effect of mutating each subset of the superpixels in the current partition. Note that $s$ is a constant independent of the size of the image. The number of steps is determined by the termination condition on the size of a single superpixel, which in the worst case is the same as a single pixel. In our setting, the algorithm terminates when a single superpixel is smaller contains fewer than 101010 pixels111rex also terminates when no combination of active superpixels has the required classification or a user-defined work budget is exceeded.. However, in general, the algorithm can continue down to the level of a single pixel, thus resulting in $n$ pixels in the last step. The algorithm performs $N$ iterations, and every iteration uses a different initial partition. The parameter $N$ is independent of the size of the image. ∎
6 Evaluation
@@
RQ4
What are suitable quality measures for explanations? A measure is suitable if it is meaningful with respect to quality and produces results that match our intuition.
@@
A causal explanation is a minimal set of pixels from an image which are sufficient to obtain the same top classification as the complete set of all pixels of an image. This immediately suggests size of explanation as a robust quality measure. As none of the tools (apart from rex) compute minimal, sufficient explanations, we use the mechanism of extraction (Section 5.3) on the output of all tools tested. We also use the same approach to compute insertion and deletion curves over all XAI outputs.
Utility and interpretation of deletion curves
(a) Starfish
(b) Bison
Figure 6: An image with a high deletion curves does not indicate a low quality explanation. The starfish (Figure 6(a)) has a normalized deletion of 0.320.320.32 because their are multiple points of high responsibility in the image. These points indicate the presence of multiple explanations. The bison (Figure 6(b)) has a low deletion of 0.10.10.1 because there is only one point of interest in the image. Deletion is a difficult measure to interpret.
When insertion and deletion curves were introduced in rise (?), the authors stated that a high insertion and a low deletion is preferable. This assertion relies on the assumption that there is only one explanation for the image’s classification, or that the XAI tool finds only one explanation. Moreover, saliency values outside of the single ‘explanation’ must be strictly uninformative: in order to keep deletion low, the saliency map cannot provide information about any substructures present in the image. Pixels must be added back into the image in an order that is essentially random. rex does not work like this: the responsibility map provides detail over the entire pixel space.
As an example, if we examine Figure 6(a), we have an image with a normalized insertion curve of 0.970.970.97 but a normalized deletion curve of 0.320.320.32. There would appear to be tension between the two numbers, as a high insertion suggests a high quality explanation whereas a high deletion suggests a low quality explanation. The mystery is resolved, however, by noting that images can and do contain multiple explanations, as first shown in (?). Indeed, there are four distinct peaks of responsibility in Figure 6(a), each of which corresponds to an independent explanation for starfish (Figure 7).
(a) Explanation 1
(b) Explanation 2
(c) Explanation 3
(d) Explanation 4
Figure 7: The responsibility map rex provides is sufficiently detailed to extract multiple valid explanations for most images. These starfish explanations were extracted from Figure 6(a)).
The deletion score is high because the responsibility map is informative even for the more lowly ranked pixels. In the multiple explanation scenario, one would expect a high deletion score. Uniquely among the tools examined here, rex supports multiple explanations for an image natively (?). In Tables 1, 3 and 5 we follow the usual procedure of indicating the lowest deletion curve value, but stress that lowest does not necessarily indicate best.
We quantify the quality of the explanations with several complementary methods. First, we calculate the minimal explanation size as a percentage of the size of the image. This coincides with Definition 3.2. It also gives an idea of how concentrated the responsibility or saliency maps provided by the XAI tools are. The smaller the area of the explanation, the more the XAI tool was able to isolate precisely those pixels (and only those pixels) required to match the original classification.
@@
Furthermore, we calculate normalized insertion and deletion curves for all images. As stated above, it is not clear how to compare the area under curve (AUC) of a low confidence image against a high confidence image to derive an average AUC performance. It is, after all, not the fault of the XAI tool that the confidence of the model is low on any particular image. To overcome this problem, we follow (?) and normalize all curves by the initial confidence of the model on image under test. Note that this can result in the the insertion curve AUC being greater than 111. This phenomenon occurs when the confidence on the overall image is lower than the confidence of intermediate insertion stages. This makes intuitive sense: if the pixel ranking is accurate, then those pixels that either do not contribute towards the classification, or reduce confidence, are left to be added last. Without these pixels, confidence is potentially much higher.
@@
For those datasets that contain segmentation information, we also measure the intersection of the minimal explanation with the segmentation. This segmentation is human-provided and we argue corresponds most closely with what a human considers acceptable. On VOC2012, for example, we would like an minimal explanation to reside mostly inside the human-provided segmentation mask and therefore contain only a small number of pixels outside the mask. For datasets that feature occluded objects, a good explanation would have as few pixels from inside the occlusion as possible, as the occlusion should not make a contribution towards the model’s classification.
@@
To answer RQ4, we argue that the size of explanations, as well as insertion curves and the intersection with the occlusion are measures that match our intuition. However, we argue that deletion curves are not, as they remain high in presence of multiple explanations.
6.1 Experimental Setup
@@
We answer research questions RQ5–RQ7 experimentally. We have implemented the proposed explanation approach in the publicly available tool rex. For all other XAI tools, apart from rise, we used the Captum PyTorch library (?) [^2] https://www.captum.ai. For rise, we used the authors’ implementation [^3] https://github.com/eclique/RISE, which we lightly altered to use PyTorch rather than tensorflow.
@@
In the evaluation, we compare rex with a wide range of explanation tools, specifically Grad-CAM (?), KernelShap (?), GradientShap (?), rise (?), lime (?), ig (?), noisetunnel (?). For the ResNet50 model only, we also include lrp. We were unable to make the captum version of lrp run on either the ViT or the ConvNext model.
While our main interest is in black-box explainability, we have included methods which require access to the model gradient, such as gradientshap and ig, and internal model layers, i.e. Grad-CAM. We call those tools which require internal access to the model ‘white-box’; those that require the access to the gradient we term ‘grey-box’, and those that just need the model’s classification, we call ‘black-box’. rex is a black-box method. PyTorch permits optional access to the gradient, so tools such as ig require at least some flexibility when calling the model. This is not guaranteed with proprietary models: hence our grey-box category.
@@
We use 444 data sets: ImageNet-1k-mini validation (?), ECSSD (?), Pascal VOC2012 (?), and a “Photobombing” dataset we created. Imagenet1k-mini comes with labels for ground truth. VOC2012 has labels and segmentation data. ECSSD comprises complex images which come with a human-provided segmentation. We created the “Photobombing” dataset by inserting black occlusions into ImageNet images, meaning we have the original label and also know the pixel coordinates of the occlusions. These occlusions are placed so as not to change the model classification. We use the torchvision [^4] https://pytorch.org/vision/stable/index.html implementations of ResNet50, ViT and ConvNext-Large with default weights on all data sets.
@@
All experiments were conducted using a server running Ubuntu 20.04 with an Nvidia A40 GPU. Moreover, all tools were used with default settings. For rex in particular, this means that it performs 20 iterations of the algorithm ($N$ in Algorithm 3), with a minimum superpixel size of 10 pixels. The pruning strategy for the work queue is “area” (the passing mutants are ordered by size) and only one item is kept in the work queue at a time, so we only refine the smallest passing mutant. Partitions are created uniformly at random. Other strategies and partitioning distributions are available in the configuration.
@@
rex uses four superpixels per partition. This is practical for images as we can split the image with just one call to a random number generator. The number of mutants produced is also relatively small (151515), which can usually be fit into a single batch for model processing. Having more initial superpixels does not lead to greater expressivity: we iteratively refine passing combinations of superpixels further, essentially recreating a more detailed initial partition produced by a greater number of superpixels. The termination conditions for the partition refinement in Algorithm 2 are: 1) the area of a superpixel is less than 101010 pixels of (resized) input image, 2) the four superpixels share the same responsibility or 3) no combination of superpixels produces the desired classification.
6.2 Results
RQ5
@@ 6.2 Results
Table 1: Results for ResNet50 model.
@@
Tables 1, 3 and 5 show the experimental results of evaluation of rex against 888 other XAI tools over 444 datasets, with three different models. The columns in the tables are: mean area (area), $\sigma$ of area (standard deviation), normalized insertion curves AUC (ins), normalized deletion curves AUC (del), proportion of explanation inside the relevant segment (IN), and the fraction of the explanation that is outside the relevant segment (OUT), shown only where appropriate. Bold indicates best result in category for all columns. Note that while we have indicated lowest AUC for deletion, lower here is not necessarily better, as we argued earlier.
@@
Table 1 shows the results with the ResNet50 model. rex consistently produces the most precise explanations. This is reinforced by the observations that rex also produces the highest insertion AUC of any of the tools. rex does particularly well on the ‘Photobombing‘ dataset, where its explanations are almost entirely (0.970.970.97) disjoint with the inserted occlusion (Figure 12). Figure 8 shows that rex has the most concentrated explanations and consistent results, with very few outliers. The grey-box tools perform worse than the other methods. Note that rex even outperforms Grad-CAM.
Table 3: Results for ViT model.
@@
Table 3 shows the results with the ‘vit_b_32‘ model from torchvision. Explanations across all tools are much larger than for ResNet50model, which seems to be a feature of the model. rex, however, still manages to produce explanations with the lowest number of redundant pixels. In general, rex, lime and Grad-CAM are the best performing tools. Even in cases where lime slightly outperforms rex (e.g., on ECSSD IN), with 0.730.730.73 of its explanation inside the segmentation against 0.720.720.72, the effect is very small. Figure 9 shows that rex again has the most concentrated explanations and consistent results, with very few outliers. lime is the closest comparable tool, though it has many more outliers and a higher median value. The significant difference in explanation size between the ResNet model and ViT model is, perhaps, of independent interest.
Table 5: Results for convnext model.
@@
Table 5 shows the results with the ‘convnext_large‘ model from torchvision. rex, lime, Grad-CAM and rise are the best performing tools. It is interesting to note that 333 of these tools fall into our ‘black-box‘ category. Figure 10 shows that rex has the best performance again, this time slightly ahead of Grad-CAM, as Grad-CAM has many more outliers than rex and a slightly larger inter-quartile range. Explanation size is comparable with that of the ResNet model, making the ViT model somewhat of an outlier.
Figure 8: Box plot of all tools over all datasets with a ResNet50 model. rex has the lowest median value and also the smallest number of outliers.
Figure 9: Box plot of all tools over all datasets with a ViT model. rex has the lowest median value and also the smallest number of outliers. ViT explanations are noticeably larger than those for ResNet50 and ConvNext.
Figure 10: Box plot of all tools over all datasets with a ConvNext model. rex has the lowest median value and also the smallest number of outliers. It even outperforms Grad-CAM, a purely white-box tool.
@@
We measured the running time of all tools. As all tools were evaluated on the same hardware with the same models and datasets, the results should be consistent. The white and gray-box methods are clearly the fastest, as expected. Indeed, Grad-CAM is essentially instantaneous, as the overhead over one call to the model is insignificant. No black-box tool can compete with this, by the nature of black-box tools. All the gradient based methods took 3$\approx$ 3≈ 3 seconds to produce initial maps. While they are still fast, the quality of their explanations is lower.
Both rex and lime took 4$\approx$ 4≈ 4s to produce maps, with rise being much slower (10$\approx$ 10≈ 10s). If we take into account the extra time for the explanation extraction (which only rex does by default), the efficiency picture changes slightly. rex’s total compute time increases only slightly (to 6$\approx$ 6≈ 6s) due the the uniformly small size of its explanations. Other tools suffer more though, even Grad-CAM’s average compute time increases to 4$\approx$ 4≈ 4s. rex is an efficient black-box tool: for RQ6 therefore, our results show that there is little trade-off between quality and compute cost.
6.3 Discussion

### v3

@@ 5.2. Iterative refinement of responsibility
OUTPUT:   a responsibility map $R:\mathds{P}_{i}\longrightarrow\mathds{Q}$
@@
7:  $R^{\prime}\leftarrow R^{\prime}\,\cup\,\mathit{iterative\_responsibility\_refinement}(x,P_{c})$
@@
One partition is rarely sufficient for a high-quality (i.e. small) explanation, therefore we compute Algorithm 1 many times and compose the results, as shown in Algorithm 2. This calculates the responsibility for each superpixel (Line 1). If the termination condition is met (Lines 2–3), the responsibility map $Q$ is updated accordingly. Otherwise, for each superpixel in $\mathds{P}_{i}$, we refine it and call the algorithm recursively. We use $\cup$ to include these newly computed values in the returned map. The algorithm terminates when: 1) the superpixels in $\mathds{P}_{i}$ are sufficiently refined (containing only very few pixels), or 2) when all superpixels in $\mathds{P}_{i}$ have the same responsibility (this condition is for efficiency). In particular, if no further subdivision of a superpixel results in the desired classification, responsibility for all superpixels is 0 and the algorithm terminates.
5.3. Explanation extraction
So far, we assume one particular partition $\mathds{P}_{i}$, which Algorithm 2 recursively refines and calculates the corresponding responsibilities of superpixels in each step by calling Algorithm 1. We note that the choice of the initial partition over the image can affect the values calculated by Algorithm 2, as this partition determines the set of possible mutants in Algorithm 1. We ameliorate the influence of the choice of any particular partition by iterating the algorithm over a set of initial partitions. Twenty iterations of the algorithm will therefore yield $20$ starting partitions, chosen at random. rex allows the user to choose from a number of different types of random to build these partitions, with uniform being the default.
@@ 5.3. Explanation extraction
4:  $R\leftarrow\mathit{iterative\_responsibility\_refinement}(x,\mathds{P}^{x})$
@@
13:  $x^{\mathit{exp}}\leftarrow$ mask pixels of $x$ that are not in $\mathcal{E}$
14:  if $\mathcal{N}(x^{\mathit{exp}})=\mathcal{N}(x)$ then
@@
The remainder of Algorithm 3 follows the method for explaining the result of an image classifier by sun2020explaining. That is, we construct a subset of pixels $\mathcal{E}$ to explain $\mathcal{N}$’s output on this particular input $x$ greedily. We add pixels to $\mathcal{E}$ as long as $\mathcal{N}$’s output on $\mathcal{E}$ does not match $\mathcal{N}(x)$. This process terminates when $\mathcal{N}$’s output is the same as on the whole image $x$. The set $\mathcal{E}$ is returned as an explanation. We note that this extraction procedure is not normally followed by other familiar XAI tools, as their several different definitions of explanation do not extend to finding minimal sets of pixels which induce the desired classification.
While we approximate the computation of an explanation in order to ensure efficiency of our approach, the algorithm is built on solid theoretical foundations, which distinguishes it from other random or heuristic-based approaches. In practice, while our algorithm uses an iterative average of a greedy approximation, it yields highly accurate results (Section 6). Furthermore, our approach is simple and general, and uses the neural network as a black-box.
5.4. Illustrative example
@@
Figure 4. The rex algorithm in action: rex creates an initial random partition of an image into 44 sections (①). All combinations of these sections are queried by the model, with further refinement applied to those sections or combinations of sections that meet the requirements (②). Some combinations, highlighted in green, are classified as bus, others, in red, are not. This results, after several iterations, in a detailed responsibility map (③), from which a minimal passing explanations can be extracted (④).
To illustrate how rex works, consider Figure 4, which is classified as 'bus' by a ResNet50, even though there is an occlusion in the middle. Initially, rex picks an arbitrary partition of the image. This results in $15$ combinations of masking superpixels, though in practice, we only need to examine $14$ combinations as the $15^{th}$ is the entire image, which we have already tested. We use Algorithm 1 to calculate the responsibility of each superpixel. Those superpixels or combinations of superpixels which contribute towards the correct classification are further refined. This iterative refinement is a recursive application of the same initial random partitioning into four superpixels, as in Algorithm 2.
The responsibility maps shown in Figure 5 demonstrate the importance of Algorithm 3. The initial partitioning of the image can greatly affect the quality of the responsibility map. Figure 5(a) shows the responsibility map after one iteration of the algorithm. While it still indicates well the areas of interest in the image it is also rather crude. Figure 5(b) contains more refined information about the distribution of responsibility over the pixels in the image. We now have two distinct peaks of responsibility on either side of the central occlusion (a person). Further iterations bring about further refinement. By iteration $100$ (Figure 5(e)), it is clear that the ‘dominant’ explanation is around the rear tire of the bus. The other explanation however, while depressed in relation to the tire, has not vanished and can still be found through searching the responsibility map (CKK25). Moreover, a disjoint explanation exists for this image, as given in Figure 6(a).
@@
Figure 5. Improvement of rex’s pixel ranking on ‘bus’ (Figure 4) as the number of iterations NN increases (Algorithm 3)
@@
Figure 6. Two different explanations extracted from the $30$ iteration ranking in Figure 5Figure 6(a) contrasts interestingly with Figure 6(b): it seems that the front tire by itself is not sufficient (a very small rectangular section of the rear tire is included in the explanation), but the rear tire alone is sufficienct.
5.5. Termination and complexity
It is clear that the algorithm necessarily terminates as soon as the parts can no longer be divided into superpixels. Moreover, the number of calls it performs to the model is linear in the size of the image, as proved in the following lemma.
@@ 5.5. Termination and complexity
The computation of responsibilities of superpixels in one partition is $O(2^{s})$, as the algorithm examines the effect of mutating each subset of the superpixels in the current partition. Note that $s$ is a constant independent of the size of the image. The number of steps is determined by the termination condition on the size of a single superpixel, which in the worst case is the same as a single pixel. In our setting, the algorithm terminates when a single superpixel is contains fewer than $10$ pixels111rex also terminates when a user-defined work budget is exceeded.. However, in general, the algorithm can continue down to the level of a single pixel, thus resulting in $n$ pixels in the last step. The algorithm performs $N$ iterations, and every iteration uses a different initial partition. The parameter $N$ is independent of the size of the image. ∎
Recall the research question RQ3:
RQ3:
Is there an efficiently computable approximation to explanations?
The complexity analysis of Algorithm 3 in Lemma 5.2 provides an affirmative answer to this research question.
6. Evaluation
@@
RQ4:
What are suitable quality measures for explanations?
@@
A causal explanation is a minimal set of pixels from an image which are sufficient to obtain the same top-$1$ classification as the complete set of all pixels of the image. This immediately suggests size of explanation as a robust quality measure. As none of the tools (apart from rex) compute minimal, sufficient explanations, we use our mechanism of extraction (Section 5.3) on the output of all tools tested.
We also use a number of complementary measures as proxies for the quality of the explanations. Insertion and deletion curves, introduced in petsiuk2018rise, are widely used to assess the quality of saliency maps. For insertion curves, the model confidence on a target class is measured as pixels are inserted over a baseline value. The order of insertion is derived from the map. If a map has accurately identified the most important pixels for a class, then the class confidence should rise quickly. This results in a large AUC (area under curve). Deletion curves are calculated in the opposite way, replacing the pixels with a baseline value. Again, if the map has identified the most important pixels accurately, then the model confidence should drop quickly, resulting in a low AUC.
@@
While a rex explanation is not a map, we calculate insertion and deletion over the responsibility map. The AUC for both insertion and deletion are obviously tied to the original confidence of the model. To allow for a fair comparison between images with different initial confidences, we follow calderon2024real and normalize all curves by the initial confidence of the model on the image under test. Note that this can result in the the insertion curve AUC being greater than $1$. This phenomenon occurs when the confidence on the entire image is lower than the confidence of intermediate insertion stages. This makes intuitive sense: if the pixel ranking is accurate, then those pixels that either do not contribute towards the classification, or reduce confidence, are left to be added last. Without these ‘negative’ pixels, confidence is potentially much higher.
@@
For those datasets that contain segmentation information, we also measure the intersection of the minimal explanation with that segmentation. This segmentation is human-provided and we argue corresponds most closely with what a human considers acceptable. On VOC2012, for example, we would like a minimal explanation to reside mostly inside the human-provided segmentation mask and therefore contain only a small number of pixels outside the mask. For datasets that feature occluded objects, a good explanation would have as few pixels from inside the occlusion region as possible, as the occlusion should not make a substantial contribution towards the model’s classification.
@@
To answer RQ4, we argue that the size of explanations, as well as insertion curves and the intersection with the human segmentation are measures that match our intuition. However, we argue that deletion curves are not, as they remain relatively high in presence of multiple detected explanations (Section 6.3).
6.1. Experimental Setup
@@
We answer research questions RQ5–RQ7 experimentally. We have implemented the proposed explanation approach in the publicly available tool rex. For all other XAI tools, apart from rise, we used the Captum PyTorch library (kokhlikyan2020captum) [^2] https://www.captum.ai. For rise, we used the authors’ implementation [^3] https://github.com/eclique/RISE, which we lightly altered to use PyTorch rather than the original TensorFlow framework.
@@
In the evaluation, we compare rex with a wide range of explanation tools, specifically Grad-CAM (CAM), KernelShap (lundberg2017unified), GradientShap (gradientshap), rise (petsiuk2018rise), lime (lime), ig (sundararajan2017axiomatic) and noisetunnel (smilkov2017smoothgradremovingnoiseadding). For the ResNet50 model only, we also include lrp (LRP). We were unable to make the captum version of lrp run on the other tested models.
While our main interest is in black-box explainability, we have included methods which require access to the model gradient, such as GradientShap and ig, and internal model layers, i.e. Grad-CAM. This is largely due to the relative paucity of purely black-box XAI methods for image data.
@@
We use $4$ data sets: ImageNet-1k-mini validation (imagenet), ECSSD (ecssd), Pascal VOC2012 (pascal-voc-2012), and a “Photobombing” dataset we created. Imagenet1k-mini comes with labels for ground truth. VOC2012 has labels and segmentation data. ECSSD comprises complex images which come with a human-provided segmentation. We created the “Photobombing” dataset by inserting black occlusions into ImageNet images, meaning we have the original label and also know the pixel coordinates of the occlusions. These occlusions are placed so as not to change the model classification. We use the torchvision [^4] https://pytorch.org/vision/stable/index.html implementations of ResNet50, ViT and ConvNext-Large with default weights on all data sets.
@@
All experiments were conducted using a server running Ubuntu 20.04 with an Nvidia A40 GPU. All tools were used with default settings. For rex in particular, this means that it performs $20$ iterations of the algorithm ($N$ in Algorithm 3), with a minimum superpixel size of $10$ pixels. The pruning strategy for the work queue is “area” (the passing mutants are ordered by size, smallest to largest) and only one item is kept in the work queue at a time, so we only refine the smallest passing mutant. Partitions are created uniformly at random. Other strategies and partitioning distributions are available in the configuration.
@@
rex uses four superpixels per partition. This is practical for images as we can split the image with just one call to a random number generator. The number of mutants produced is also relatively small ($15$), which can usually be fit into a single batch for model inference. Having more initial superpixels does not lead to greater expressivity: we iteratively refine passing combinations of superpixels, essentially recreating the more detailed initial partition which would be produced by a greater number of superpixels. The termination conditions for the partition refinement in Algorithm 2 are: 1) the area of a superpixel is less than $10$ pixels of (resized) input image, or 2) the four superpixels share the same responsibility.
6.2. Results
RQ5:
@@ 6.2. Results
Table 1. Results for ResNet50 model.
@@
Tables 1, 3 and 5 show the experimental results of evaluation of rex against $8$ other XAI tools over $4$ datasets, with three different models. The columns in the tables are: mean area (area), $\sigma$ of area (standard deviation), normalized insertion curve AUC (ins), normalized deletion curve AUC (del), proportion of explanation inside the relevant segment (IN), and the fraction of the explanation that is outside the relevant segment (OUT), shown only where appropriate. Bold indicates best result in category for all columns. Note that while we have indicated lowest AUC for deletion as ‘best’, lower here is not necessarily better, as we argue in Section 6.3.
@@
Table 1 shows the results with the ResNet50 model. rex consistently produces the most smallest explanations. This is reinforced by the observations that rex also produces the highest insertion AUC of any of the tools. rex does particularly well on the ‘Photobombing‘ dataset, where its explanations are almost entirely ($0.97$) disjoint from the inserted occlusion (Figure 13). Figure 7 shows that rex has the most concentrated explanations and consistent results, with very few outliers. The grey-box tools perform worse than the other methods. Note that rex even outperforms Grad-CAM, a white-box method.
Table 3. Results for ViT model.
@@
Table 3 shows the results with the ‘vit_b_32’ model from torchvision. Explanations across all tools are much larger than for the ResNet50, which seems to be a feature of this model. rex, however, still manages to produce explanations with the lowest number of redundant pixels. In general, rex, lime and Grad-CAM are the best performing tools. Even in cases where lime slightly outperforms rex (e.g., on ECSSD IN), with $0.73$ of its explanation inside the segmentation against $0.72$, the effect is very small. Figure 8 shows that rex again has the most concentrated explanations and consistent results, with very few outliers. lime is the closest comparable tool, though it has many more outliers and a higher median value. The significant difference in explanation size between the ResNet model and ViT model is, perhaps, of independent interest.
Table 5. Results for convnext model.
@@
Table 5 shows the results with the ‘convnext_large’ model from torchvision. rex, lime, Grad-CAM and rise are the best performing tools. It is interesting to note that $3$ of these tools fall into our black-box category. Figure 9 shows that rex has the best performance again, this time slightly ahead of Grad-CAM, as Grad-CAM has many more outliers than rex and a slightly larger inter-quartile range. Explanation size is comparable with that of the ResNet model, revealing the ViT model to be somewhat of an outlier.
Figure 7. Box plot of all tools over all datasets with a ResNet50 model. rex has the lowest median value and also the smallest number of outliers.
Figure 8. Box plot of all tools over all datasets with a ViT model. rex has the lowest median value and also the smallest number of outliers. ViT explanations are noticeably larger than those for ResNet50 and ConvNext.
Figure 9. Box plot of all tools over all datasets with a ConvNext model. rex has the lowest median value and also the smallest number of outliers. It even outperforms Grad-CAM, a purely white-box tool.
@@
We measured the running time of all tools. As all tools were evaluated on the same hardware with the same models and datasets, the results should be consistent. The white and grey-box methods are clearly the fastest, as expected. Indeed, Grad-CAM is essentially instantaneous, as the overhead over one call to the model is insignificant. No black-box tool can compete with this, by the nature of black-box tools. All the gradient-based methods took $\approx 3$ seconds to produce initial maps. Note that this is the time taken to generate initial saliency, minimal pixel subsets still need to be extracted.
Both rex and lime took $\approx 4$s to produce maps, with rise being much slower ($\approx 10$s). Taking into account the extra time for the explanation extraction (which only rex does by default), the efficiency picture changes slightly. rex’s total compute time increases only slightly (to $\approx 6$s) due the the uniformly small size of its explanations. Other tools suffer more, even Grad-CAM’s average compute time increases to a total of $\approx 4$s. This is due to the relatively noisy output of Grad-CAM, requiring more work from the extraction algorithm as a result. rex is an efficient black-box tool: for RQ6 therefore, our results show that there is little trade-off between quality and compute cost.
6.3. Discussion
(a) Starfish
(b) Bison
Figure 10. An image with a high deletion curves does not indicate a low quality explanation. The starfish (Figure 10(a)) has a normalized deletion of $0.32$ because their are multiple points of high responsibility in the image. These points indicate the presence of multiple explanations. The bison (Figure 10(b)) has a low deletion of $0.1$ because there is only one point of interest in the image. Deletion is a difficult measure to interpret.
Utility and interpretation of deletion curves
petsiuk2018rise stated that a high insertion and a low deletion is preferable. This assertion relies on the assumption that there is only one explanation for the image’s classification, or that the XAI tool finds only one explanation. For deletion AUC to remain low, saliency values outside of the single ‘explanation’ must be strictly uninformative: the saliency map cannot provide information about any other substructures present in the image. It must destroy that signal. Essentially, pixel deletion after the first discovered explanatory pixel set must be close to random, breaking up information from other potential explanations. Given that a large body of work is dedicated to ‘cleaning up’ saliency maps, especially from Grad-CAM-style tools (noisetunnel; smilkov2017smoothgradremovingnoiseadding; gradcamplusplus), it is not surprising that this genuine signal is lost.
rex does not work like this: the responsibility map provides detailed information over the entire pixel space. Because the map is not itself the explanation, there is no motivation to remove or delete ‘inconvenient’ hot-spots. Moreover, because rex is not using activation, which can be noisy, every hot-spot in the rex map is indeed associated to some degree with the desired classification. rex utilizes the richness of its responsibility map to discover multiple, different explanations of an image (CKK25). This is unlike the other tools in our experimental comparison, all of which return one explanation by design.
As an example, if we examine Figure 10(a), we have an image with a normalized insertion curve of $0.97$ but a normalized deletion curve of $0.32$. There would appear to be tension between the two numbers, as a high insertion suggests a high quality explanation whereas a high deletion suggests a low quality explanation. The mystery is resolved, however, by noting that the image contains multiple sufficient explanations. Indeed, there are four distinct peaks of responsibility in Figure 10(a), each of which corresponds to an independent explanation for starfish (Figure 11). It takes a long time to delete all of the pixels from the well-defined peaks in the responsibility map.
(a) Explanation 1
(b) Explanation 2
(c) Explanation 3
(d) Explanation 4
Figure 11. The responsibility map rex provides is sufficiently detailed to extract multiple valid explanations for most images. These starfish explanations were extracted from Figure 10(a).
In Tables 1, 3 and 5 we follow the usual procedure of indicating the lowest deletion curve value, but stress that lowest does not necessarily indicate best.
Differences between methods